# Hallo3: Highly Dynamic and Realistic Portrait Image Animation with Diffusion Transformer Networks

### v2

@@ 4.1 Experimental Setups
Implementation. We initialize the identity reference and denoising networks with weights derived from CogVideoX-5B-I2V[41]. During both training phases, we employ the v-prediction diffusion loss[27] for optimization. Each training phase comprises 20,000 steps, utilizing 64 NVIDIA H100 GPUs. The batch size per GPU is set to 1, with a learning rate of $480\times 720$ pixels. To enhance video generation variability, the reference image, guidance audio and textual prompt are dropped with a probability of 0.05 during training.
Evaluation Metrics. We employed a range of evaluation metrics for generated videos across benchmark datasets, including HDTF [47] and Celeb-V [50]. These metrics comprise Fréchet Inception Distance (FID) [29], Fréchet Video Distance (FVD) [36], Synchronization-C (Sync-C) [6], Synchronization-D (Sync-D) [6], and E-FID [35]. FID and FVD quantify the similarity between generated images and real data, while Sync-C and Sync-D assess lip synchronization accuracy. E-FID evaluates the image quality based on features extracted from the Inception network. Besides, we introduced VBench [13] metrics to enhance evaluation, focusing on dynamic degree and subject consistency. Dynamic degree is measured using RAFT [34] to quantify the extent of motion in generated videos, providing a comprehensive assessment of temporal quality. Subject consistency is measured through DINO [5] feature similarity, ensuring uniformity of a subject’s appearance across frames.
Baseline Approaches. We considered several representative audio-driven talking face generation methods for comparison, all of which have publicly available source code or implementations. These methods include SadTalker [45], DreamTalk [19], AniPortrait [38], and Hallo [39, 8]. The selected approaches encompass both GANs and diffusion models, as well as techniques utilizing intermediate facial representations alongside end-to-end frameworks. This diversity in methodologies allows for a comprehensive evaluation of the effectiveness of our proposed approach compared to existing solutions.
@@ 4.2 Comparison with State-of-the-art
Comparison on HDTF and Celeb-V Dataset. As shown in Table 3 and  3, our method achieves best results on FID, FVD on both datasets. Although our approach shows some disparity compared to the state-of-the-art in lip synchronization, it still demonstrates promising results as illustrated in Figure 6. This is because, to generate animated portraits from different perspectives, our training data primarily consists of talking videos with significant head and body movements, as well as diverse dynamic scenes, unlike static scenes with minimal motion. While this may lead to some performance degradation on lip synchronization, it better reflects realistic application scenarios.
@@ 4.4 Generation Controllability
Textual Prompt for Subject Animation. To evaluate whether textual conditional controllability is effectively preserved, we conducted a series of experiments comparing the performance of our method to that of the baseline model, CogVideoX [41], using same text prompts. As shown in Figure 11, the results shows that our model maintains its ability for textual control, and effectively captures the interaction between different subjects as dictated by the textual prompts.
@@ 5 Conclusion
This paper introduces advancements in portrait image animation utilizing the enhanced capabilities of a transformer-based diffusion model. By integrating audio conditioning through cross-attention mechanisms, our approach effectively captures the intricate relationship between audio signals and facial expressions, achieving substantial lip synchronization. To preserve facial identity across video sequences, we incorporate an identity reference network. Additionally, we utilize motion frames to enable the model to generate long-duration video extrapolations. Our model produces animated portraits from diverse perspectives, seamlessly blending dynamic foreground and background elements while maintaining temporal consistency and high fidelity.

### v4

@@ 4.1 Experimental Setups
Implementation. We initialize the identity reference and denoising networks with weights derived from CogVideoX-5B-I2V[yang2024cogvideox]. During both training phases, we employ the v-prediction diffusion loss[salimans2022progressive] for optimization. Each training phase comprises 20,000 steps, utilizing 64 NVIDIA H100 GPUs. The batch size per GPU is set to 1, with a learning rate of $1\times 10^{-5}$. The resolution of the training videos is 480 x 720 pixels. To enhance video generation variability, the reference image, guidance audio, and textual prompt are dropped with a probability of 0.05 during training.
Evaluation Metrics. We employed a range of evaluation metrics for generated videos across benchmark datasets, including HDTF and Celeb-V. These metrics comprise Fréchet Inception Distance (FID) [Seitzer2020FID], Fréchet Video Distance (FVD) [unterthiner2018towards], Synchronization-C (Sync-C) [Chung16a], Synchronization-D (Sync-D) [Chung16a], and E-FID [tian2024emo]. FID and FVD quantify the similarity between generated images and real data, while Sync-C and Sync-D assess lip synchronization accuracy. E-FID evaluates image quality based on features extracted from the Inception network. Besides, we introduced V-bench [huang2023vbench] metrics to enhance evaluation, focusing on dynamic degree and subject consistency. Dynamic degree is measured using RAFT [teed2020raft] to quantify the extent of motion in generated videos, providing a comprehensive assessment of temporal quality. Subject consistency is measured through DINO feature similarity, ensuring uniformity of a subject’s appearance across frames.
Baseline Approaches. We considered several representative audio-driven talking face generation methods for comparison, all of which have publicly available source code or implementations. These methods include SadTalker [zhang2022sadtalker], DreamTalk [ma2023dreamtalk], AniPortrait [wei2024aniportrait], and Hallo [xu2024hallo, cui2024hallo2]. The selected approaches encompass both GANs and diffusion models, as well as techniques utilizing intermediate facial representations alongside end-to-end frameworks. This diversity in methodologies allows for a comprehensive evaluation of the effectiveness of our proposed approach in comparison to existing solutions.
@@ 4.2 Comparison with State-of-the-art
Comparison on HDTF and Celeb-V Dataset. As shown in Table 3 and  3, our method achieves best results on FID, FVD on both datasets. Although our approach shows some disparity compared to the state-of-the-art (SOTA) in lip synchronization, it still demonstrates promising results as illustrated in Figure 6. This is because, to generate animated portraits from different perspectives, our training data primarily consists of talking videos with significant head and body movements, as well as diverse dynamic scenes, unlike static scenes with minimal motion. While this may lead to some performance degradation on lip synchronization, it better reflects realistic application scenarios.
@@ 4.4 Generation Controllability
Textual Prompt for Subject Animation. To evaluate whether textual conditional controllability is effectively preserved, we conducted a series of experiments comparing the performance of our method to that of the baseline model, CogVideoX [yang2024cogvideox], using same text prompts. As shown in Figure 11, the results shows that our model maintains its ability for textual control, and effectively captures the interaction between different subjects as dictated by the textual prompts.
@@ 5 Conclusion
This paper introduces advancements in portrait image animation utilizing the enhanced capabilities of a DiT-based diffusion model. By integrating audio conditioning through cross-attention mechanisms, our approach effectively captures the intricate relationship between audio signals and facial expressions, achieving substantial lip synchronization. To preserve facial identity across video sequences, we incorporate an identity reference network. Additionally, we utilize motion frames to enable the model to generate long-duration video extrapolations. Our model produces animated portraits from diverse perspectives, seamlessly blending dynamic foreground and background elements while maintaining temporal consistency and high fidelity.